# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ Appendix B Example: Insecurity of Interferometric BB84
In this appendix, we show how to employ the Reversed-Space method on an interferometric BB84 scheme to find the attack described in Section 5.3. In interferometric BB84 schemes, two possible times and a single photon are used to encode a qubit. One basis can be defined by the photon entering at one time or the other, while the other basis is comprised of a superposition of the two times with different phase values. When only states that use both times in a superposition are used, the scheme is called time-multiplexed, phase-encoded. Such schemes were initially suggested by [12] and implemented by Townsend [79] and many others (e.g., [80, 81, 82, 83, 84, 33], see as well [85, 5]. Encoding and multiplexing methods are further discussed in Section 8).
In order to produce and measure such superpositioned pulses, it is common to use interferometers (see below and Appendix A). Yet, once a protocol is implemented via photons and interferometers, two immediate reasons cause an enlargement of the quantum space in use: [^5] There are other possible causes for space enlargement. For instance, the shape of the pulse (in the frequency domain and/or the time domain) is not generated by Alice or measured by Bob in an ideal way, and this opens another source space enlargement. first, interferometers inherently introduce a higher-dimension space; and second, having pulses with zero photons, or more than one photon, implies a higher dimension as well.
In this subsection, we demonstrate a Reversed-Space Attack on an interferometric BB84 implementation. The attack we define here exposes a vulnerability inherent to a large class of implementations, and is directly applicable to the implementations in [86, 87, 88] and the NICT-NEC implementation in [89].
Our focus in this subsection is on a single physical apparatus. Note that not all interferometric QKD implementations are insecure in this fashion [28], but other implementation methods should still be analyzed [90, 91].
@@
B.1 Interferometric Implementation of BB84
Consider a BB84 implementation which uses two time-separated modes (pulses). For every transmission, the first mode arrives to Bob’s lab at time $t_{0}^{\prime}$, and the second mode at $t_{1}^{\prime}=t_{0}^{\prime}+\Delta T$. We denote these pulses as $|{t^{\prime}_{0}}\rangle$ and $|{t^{\prime}_{1}}\rangle$ respectively. The users use the Hadamard and computational bases, so that an ideal Alice sends one of the following four states,
$|{0}\rangle_{\mathrm{A}}\equiv|{t^{\prime}_{0}}\rangle$ $|{+}\rangle_{\mathrm{A}}\equiv\left(|{t^{\prime}_{0}}\rangle+|{t^{\prime}_{1}}\rangle\right)/\sqrt{2}\phantom{\text{ .}}$
@@ B.1 Interferometric Implementation of BB84
Bob measures the qubit using a Mach-Zehnder interferometer, which is a device composed of two beam splitters (BS) with one short path, one long path, and a controlled phase shifter $P_{\phi}$, that is placed at the long arm of the interferometer. (See Appendix A for a mode details on linear-optics devices, and analysis of their operation on photonic states). The length difference between the two arms is determined by $\Delta T$: when the first pulse travels through the long arm, and the second through the short arm, they arrive together at the output. Due to that exact timing of the pulses, each incoming qubit is transformed into a superposition of 6 possible modes: 3 time modes ($t_{0}$, $t_{1}$, $t_{2}$) at the straight ($s$) output arm of the interferometer, and 3 modes at the down ($d$) output arm; see Figure 6.
[…]
Figure 6: A Mach-Zehnder interferometer. (a) An input qubit. The time-difference between the two incoming modes is identical to the difference between the two arms; (b) a vacuum state entering the second (blocked) arm; (c) beam splitters; (d) phase shifter $P_{\phi}$; (e) six output modes.
For the sake of simplicity we denote these modes as $s_{0},s_{1},s_{2},d_{0},d_{1},d_{2}$, and since we only consider pulses with zero or one photon, we can use the states, $|{s_{0}}\rangle$, $|{d_{0}}\rangle$, etc. [^6] Using the Fock-space notations (Section 2.2) and the description of interferometers in Appendix A, a basis state in Bob’s space is $|{n_{s_{0}},n_{s_{1}},n_{s_{2}},n_{d_{0}},n_{d_{1}},n_{d_{2}}}\rangle^{\!\text{\tiny F}}$, and we define $|{100000}\rangle^{\!\text{\tiny F}}\equiv|{s_{0}}\rangle$; $|{010000}\rangle^{\!\text{\tiny F}}\equiv|{s_{1}}\rangle$; $|{001000}\rangle^{\!\text{\tiny F}}\equiv|{s_{2}}\rangle$; $|{000100}\rangle^{\!\text{\tiny F}}\equiv|{d_{0}}\rangle$; $|{000010}\rangle^{\!\text{\tiny F}}\equiv|{d_{1}}\rangle$; $|{000001}\rangle^{\!\text{\tiny F}}\equiv|{d_{2}}\rangle$, and the vacuum state $|{000000}\rangle^{\!\text{\tiny F}}\equiv|{V}\rangle$., along with the vacuum state $|{V}\rangle$ (a pulse that contains no photons in any of the modes).
The interferometer evolves these states according to $|{V}\rangle\mapsto|{V}\rangle_{\mathrm{B}}$ and
$\begin{split}|{t^{\prime}_{0}}\rangle&\mapsto(|{s_{0}}\rangle_{\mathrm{B}}-e^{i\phi}|{s_{1}}\rangle_{\mathrm{B}}+i|{d_{0}}\rangle_{\mathrm{B}}+ie^{i\phi}|{d_{1}}\rangle_{\mathrm{B}})/2\\ |{t^{\prime}_{1}}\rangle&\mapsto(|{s_{1}}\rangle_{\mathrm{B}}-e^{i\phi}|{s_{2}}\rangle_{\mathrm{B}}+i|{d_{1}}\rangle_{\mathrm{B}}+ie^{i\phi}|{d_{2}}\rangle_{\mathrm{B}})/2.\end{split}$ (25)
In our example, Bob fixes the phase $\phi$ to 0. Thus, Alice’s qubit evolves in the interferometer as
@@
B.2 Identifying the Reversed Space of the Interferometric Setup
@@
We now follow the framework defined in the last subsections and specify Bob’s operation and the space he measures, in two simple cases. We then derive the corresponding reversed space that applies in each setting.
Example: when Bob measures two modes.
Assume that Bob only measures the two modes that correspond to time $t_{1}$, namely, $|{s_{1}}\rangle$ and $|{d_{1}}\rangle$. [^7] This setting happens, for instance, when the BB84 protocol is run using the Hadamard and $y$ bases.
Thus, the space Bob measures, $\mathcal{H}^{\mathrm{B}}$, is spanned by $B=\{|{V}\rangle,|{s_{1}}\rangle,|{d_{1}}\rangle\}$. We can reverse each mode by evolving it backwards in time through the interferometer. As shown in Appendix A, a reversal of a single mode through the interferometer is given by
$\begin{split}|{s_{n}}\rangle&\mapsto(\phantom{-i}|{a_{n}}\rangle-\phantom{i}e^{-i\phi}|{a_{n-1}}\rangle-i|{b_{n}}\rangle-ie^{-i\phi}|{b_{n-1}}\rangle)/2\\ |{d_{n}}\rangle&\mapsto(-i|{a_{n}}\rangle-ie^{-i\phi}|{a_{n-1}}\rangle-\phantom{i}|{b_{n}}\rangle+\phantom{i}e^{-i\phi}|{b_{n-1}}\rangle)/2.\end{split}$ (27)
Where $|{a_{n}}\rangle$ is a pulse in the input arm at time $n$, and $|{b_{n}}\rangle$ is a pulse in the blocked arm at time $n$. The derivation of the formula is given in Appendix A.
Reversing $|{s_{1}}\rangle$ and $|{d_{1}}\rangle$ and then “tracing out” the blocked ancillary system (the ‘$b$’ arm) implies that $\mathcal{H}^{\mathrm{P}}$ must include at least the span of $\{|{V}\rangle,|{t^{\prime}_{1}}\rangle,|{t^{\prime}_{0}}\rangle\}$.
Example: when Bob measures six modes.
@@
Now suppose that Bob measures time-bins $t_{0},t_{1},t_{2}$, so his measured space $\mathcal{H}^{\mathrm{B}}$ is the span of $\{|{V}\rangle$, $|{d_{0}}\rangle$, $|{d_{1}}\rangle$, $|{d_{2}}\rangle$, $|{s_{0}}\rangle$, $|{s_{1}}\rangle$, $|{s_{2}}\rangle\}$. Again using the reversed transformation in Eq. (27), we get that the reversed space $\mathcal{H}^{\mathrm{P}}$ is much larger than $\mathcal{H}^{\mathrm{A}}$: it is the space that allows the photon to be in any superposition of time modes $t^{\prime}_{-1}$ to $t^{\prime}_{2}$. This means that the analysis needs only focus on the space spanned by $\{|{V}\rangle$, $|{t^{\prime}_{-1}}\rangle$, $|{t^{\prime}_{0}}\rangle$, $|{t^{\prime}_{1}}\rangle$, $|{t^{\prime}_{2}}\rangle\}$.
B.3 Attack on BB84 Using the Hadamard and Computational Bases: When Bob Measures Six Modes
We will now analyze the interferometric setting described in Section B.1. After defining the interesting spaces for this setting (as shown in the example above), we now try to find oblivious attacks, using the formalization of Section 5.2.3. First note that Bob’s unitary, $U_{\mathrm{B}}$, is the same for both the computational and the Hadamard bases, $\beta^{C}=\beta^{H}$, and is characterized by
@@ B.3 Attack on BB84 Using the Hadamard and Computational Bases: When Bob Measures Six Modes
$\beta^{H}_{\begin{array}[t]{l}\scriptstyle k=\{t^{\prime}_{-1},t^{\prime}_{0},t^{\prime}_{1},t^{\prime}_{2}\},\\[-3.0pt] \scriptstyle j=\{s_{0},s_{1},s_{2},d_{0},d_{1},d_{2}\}\end{array}}\!\!\!\!\!\!=\frac{1}{2}\left(\begin{array}[]{cccccc}-1&0&0&i&0&0\\ 1&-1&0&i&i&0\\ 0&1&-1&0&i&i\\ 0&0&1&0&0&i\end{array}\right),$
which is immediately given by extending Eq. (25) with $\phi=0$ to times $t^{\prime}_{-1}$ and $t^{\prime}_{2}$.
@@
Denote with $B=\{|{V}\rangle$, $|{d_{0}}\rangle$, $|{d_{1}}\rangle$, $|{d_{2}}\rangle$, $|{s_{0}}\rangle$, $|{s_{1}}\rangle$, $|{s_{2}}\rangle\}$ a basis of Bob’s measured space $\mathcal{H}^{\mathrm{B}}$. When Bob measures in the Hadamard basis, he interprets his measurement in the following way, $J_{0}=\{|{d_{1}}\rangle\}$; $J_{1}=\{|{s_{1}}\rangle\}$; $J_{\operatorname{loss}}=B\setminus(J_{0}\cup J_{1})$; and $J_{\operatorname{invalid}}=\emptyset$. [^8] We limit the analysis to pulses that contain at most a single photon. Under this assumption, there are no invalid states for this setting. Consider the case where Alice sends $|{+}\rangle$, namely, $\alpha_{t^{\prime}_{0}}=\alpha_{t^{\prime}_{1}}=\frac{1}{\sqrt{2}}$. An error occurs if Bob measures $|{s_{1}}\rangle$, $J_{\operatorname{error}}=\{|{s_{1}}\rangle\}$, and by Observation 1, our attack is required to satisfy
$\frac{1}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{0}}|{E_{t^{\prime}_{0},t^{\prime}_{0}}}\rangle_{\mathrm{E}}+\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle_{\mathrm{E}})+\frac{1}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle_{\mathrm{E}}+\epsilon_{t^{\prime}_{1},t^{\prime}_{1}}|{E_{t^{\prime}_{1},t^{\prime}_{1}}}\rangle_{\mathrm{E}})=0.$ (28)
@@
$\frac{i}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{0}}|{E_{t^{\prime}_{0},t^{\prime}_{0}}}\rangle_{\mathrm{E}}-\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle_{\mathrm{E}})+\frac{i}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle_{\mathrm{E}}-\epsilon_{t^{\prime}_{1},t^{\prime}_{1}}|{E_{t^{\prime}_{1},t^{\prime}_{1}}}\rangle_{\mathrm{E}})=0\text{.}$ (29)
@@
$\begin{split}i\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle+i\epsilon_{t^{\prime}_{0},t^{\prime}_{2}}|{E_{t^{\prime}_{0},t^{\prime}_{2}}}\rangle=0,\quad\quad\quad-\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle+\epsilon_{t^{\prime}_{0},t^{\prime}_{2}}|{E_{t^{\prime}_{0},t^{\prime}_{2}}}\rangle=0,\end{split}$ (30)
@@
$\begin{split}i\epsilon_{t^{\prime}_{1},t^{\prime}_{-1}}|{E_{t^{\prime}_{1},t^{\prime}_{-1}}}\rangle+i\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle=0,\quad\quad\quad-\epsilon_{t^{\prime}_{1},t^{\prime}_{-1}}|{E_{t^{\prime}_{1},t^{\prime}_{-1}}}\rangle+\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle=0,\end{split}$ (31)
@@
$\begin{split}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}&\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}p|{\phi}\rangle_{\mathrm{E}}|{t^{\prime}_{0}}\rangle_{\mathrm{P}}+p_{1}|{\phi_{1}}\rangle|{t^{\prime}_{-1}}\rangle_{\mathrm{P}}+p_{2}|{\psi_{0}}\rangle_{\mathrm{E}}|{V}\rangle_{\mathrm{P}},\\ |{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}&\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}p|{\phi}\rangle_{\mathrm{E}}|{t^{\prime}_{1}}\rangle_{\mathrm{P}}+p_{3}|{\phi_{2}}\rangle|{t^{\prime}_{2}}\rangle_{\mathrm{P}}+p_{4}|{\psi_{1}}\rangle_{\mathrm{E}}|{V}\rangle_{\mathrm{P}},\end{split}$ (32)
@@
with orthogonal $|{E_{1}}\rangle$, $|{E_{2}}\rangle$. Readers might notice that this attack is similar to “faked states” attacks [19, 20] described in Section 3.2.3; we explore the connection between the two in Section 8.

### v3

@@ Appendix B Example: Insecurity of Interferometric BB84
In this appendix, we show how to employ the Reversed-Space method on an interferometric BB84 scheme and obtain the attack mentioned in Section 5.3.
In interferometric BB84 schemes, a qubit is encoded using a single photon in two possible time-bins. One basis can be defined by the photon arriving at one time or the other. Other qubit states are superpositions of these two times. When only states that use both times in a superposition are used, the scheme is called time-multiplexed, phase-encoded. Such schemes were implemented by Townsend (Townsend1994, ) and many others (e.g., (Hughes2000, ; Kimura2004, ; Gobby2004, ; Yuan2009, ; Lucamarini2013, ; Yuan2018, ); see also (Gisin2002, ; Xu2020, )).
In order to produce and measure pulses with time superposition, it is common to use interferometers (see below and Appendix A). Yet, once a protocol is implemented using photons and interferometers, there are two immediate reasons for an expansion of the quantum space in use. First, interferometers inherently introduce a higher-dimensional space. Second, having pulses with zero or more than one photon also implies a higher dimension. [^5] There are also other possible causes for space enlargement. For example, this could happen due to imperfect generation or measurement of the pulse shape in either the frequency or time domain.
In this appendix, we demonstrate a Reversed-Space Attack on an interferometric BB84 implementation (described below). The attack we define here exposes a vulnerability inherent to a large class of implementations, and is directly applicable to the implementations in (Walton2003, ; Nambu2004, ; Jaeger2006, ) and the NICT-NEC implementation in (Sasaki2011, ).
Our focus in this appendix is on a specific physical apparatus. Note that not all physical implementations of interferometric QKD are insecure in this fashion (gelles_security_2012, ), but other implementation methods should still be analyzed (Boaron2018, ; Beutel2021, ).
@@
B.1. Interferometric Implementation of BB84
Consider an implementation of BB84 that uses two time-separated modes (pulses) to encode a qubit. For every transmission, the first mode arrives to Bob’s lab at time $t_{0}^{\prime}$, and the second mode at $t_{1}^{\prime}=t_{0}^{\prime}+\Delta T$. We denote these pulses as $|{t^{\prime}_{0}}\rangle$ and $|{t^{\prime}_{1}}\rangle$, respectively. The ideal Alice sends one of the following four states,
$|{0}\rangle_{\mathrm{A}}\equiv|{t^{\prime}_{0}}\rangle$ $|{+}\rangle_{\mathrm{A}}\equiv\left(|{t^{\prime}_{0}}\rangle+|{t^{\prime}_{1}}\rangle\right)/\sqrt{2}\phantom{\text{ ,}}$
@@ B.1. Interferometric Implementation of BB84
where $\{|{0}\rangle_{\mathrm{A}},|{1}\rangle_{\mathrm{A}}\}$ is the computational basis, and $\{|{+}\rangle_{\mathrm{A}},|{-}\rangle_{\mathrm{A}}\}$ is the Hadamard basis.
Bob measures the qubit using a Mach-Zehnder interferometer, which is a device composed of two beam splitters (BS) with one short path, one long path, and a controlled phase shifter $P_{\phi}$, that is placed at the long arm of the interferometer. (See Appendix A for details on linear-optics devices and an analysis of their operation on photonic states). The length difference between the two arms is determined by $\Delta T$: when the first pulse travels through the long arm, and the second through the short arm, they arrive together at the output. Due to that exact timing of the pulses, each incoming qubit is transformed into a superposition of six possible modes: three time modes ($t_{0}$, $t_{1}$, $t_{2}$) at the straight ($s$) output arm of the interferometer, and three modes at the down ($d$) output arm; see Figure 4.
For simplicity, we denote these modes as $s_{0},s_{1},s_{2},d_{0},d_{1},d_{2}$. In order to construct this attack, it is sufficient to only consider photons with zero or one photon. Thus, we use the states we can use the states $\{|{s_{j}}\rangle$, $|{d_{j}}\rangle:j\in\{0,1,2\}\}$ along with the vacuum state $|{V}\rangle$ which denotes a pulse that has no photons in any of the modes. [^6] Using the Fock-space notations (Section 2.2) and the description of interferometers in Appendix A, a basis state in Bob’s space is $|{n_{d_{2}},n_{d_{1}},n_{d_{0}},n_{s_{2}},n_{s_{1}},n_{s_{0}}}\rangle^{\!\text{\tiny F}}$, and we define $|{0{,}0{,}0{,}0{,}0{,}1}\rangle^{\!\text{\tiny F}}\equiv|{s_{0}}\rangle$; $|{0{,}0{,}0{,}0{,}1{,}0}\rangle^{\!\text{\tiny F}}\equiv|{s_{1}}\rangle$; $|{0{,}0{,}0{,}1{,}0{,}0}\rangle^{\!\text{\tiny F}}\equiv|{s_{2}}\rangle$; $|{0{,}0{,}1{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\equiv|{d_{0}}\rangle$; $|{0{,}1{,}0{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\equiv|{d_{1}}\rangle$; $|{1{,}0{,}0{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\equiv|{d_{2}}\rangle$, and the vacuum state $|{0{,}0{,}0{,}0{,}0{,}0}\rangle^{\!\text{\tiny F}}\equiv|{V}\rangle$.
As shown in Appendix A interferometer evolves these states according to $|{V}\rangle\mapsto|{V}\rangle_{\mathrm{B}}$ and
(25) $\begin{split}|{t^{\prime}_{0}}\rangle&\mapsto(|{s_{0}}\rangle_{\mathrm{B}}-e^{i\phi}|{s_{1}}\rangle_{\mathrm{B}}+i|{d_{0}}\rangle_{\mathrm{B}}+ie^{i\phi}|{d_{1}}\rangle_{\mathrm{B}})/2\\ |{t^{\prime}_{1}}\rangle&\mapsto(|{s_{1}}\rangle_{\mathrm{B}}-e^{i\phi}|{s_{2}}\rangle_{\mathrm{B}}+i|{d_{1}}\rangle_{\mathrm{B}}+ie^{i\phi}|{d_{2}}\rangle_{\mathrm{B}})/2.\end{split}$
In our example, Bob fixes the phase $\phi$ to 0, regardless of the basis in which he wishes to measure; thus, $U_{\mathrm{B}}$ is fixed and identical to the operation of the interferometer. Thus, Alice’s qubit evolves in the interferometer as
@@
B.2. Identifying the Reversed Space of the Interferometric Setup
@@
We now follow the framework defined in Section 5 and specify the space Bob measures and its reversed space. We then derive the corresponding reversed space that applies in each setting.
Let us first analyze what states affect a single mode measured by Bob. As derived in Appendix A (Eq. (24)), a reversal of a single mode through the interferometer (i.e., $U_{\mathrm{B}}^{\dagger}$) is given by
(27) $\begin{split}|{s_{n}}\rangle&\mapsto(\phantom{-i}|{a_{n}}\rangle-\phantom{i}e^{-i\phi}|{a_{n-1}}\rangle-i|{b_{n}}\rangle-ie^{-i\phi}|{b_{n-1}}\rangle)/2\\ |{d_{n}}\rangle&\mapsto(-i|{a_{n}}\rangle-ie^{-i\phi}|{a_{n-1}}\rangle-\phantom{i}|{b_{n}}\rangle+\phantom{i}e^{-i\phi}|{b_{n-1}}\rangle)/2\text{,}\end{split}$
where $|{a_{n}}\rangle$ is a pulse in the input arm of the interferometer at time $t_{n}$, and $|{b_{n}}\rangle$ is a pulse in the blocked arm of the interferometer at time $t_{n}$.
@@
Now consider the six modes that Bob potentially measures in every BB84 transmission, i.e., time-bins $t_{0},t_{1},t_{2}$ in both output arms. Bob’s measured space $\mathcal{H}^{\mathrm{B}}$ is the span of $\{|{V}\rangle$, $|{d_{0}}\rangle$, $|{d_{1}}\rangle$, $|{d_{2}}\rangle$, $|{s_{0}}\rangle$, $|{s_{1}}\rangle$, $|{s_{2}}\rangle\}$. We now use the reversed transformation in Eq. (27) to derive the reversed space $\mathcal{H}^{\mathrm{P}}$. After “tracing out” the blocked ancillary system (the “bb” arm), we get that $\mathcal{H}^{\mathrm{P}}$ is the space that allows the photon to be in any superposition of time modes $t^{\prime}_{-1}$ to $t^{\prime}_{2}$ of the interferometer input (the “aa” arm). Surprisingly, this space is much larger than $\mathcal{H}^{\mathrm{A}}$.
Now, our analysis can only focus on the space spanned by $\{|{V}\rangle$, $|{t^{\prime}_{-1}}\rangle$, $|{t^{\prime}_{0}}\rangle$, $|{t^{\prime}_{1}}\rangle$, $|{t^{\prime}_{2}}\rangle\}$.
B.3. A Reversed-Space Attack
We now design an oblivious attack on the BB84 realization described in Section B.1, using the reversed-space methodology described in Section 5.2.3.
As mentioned above, Bob’s unitary $U_{\mathrm{B}}$ is the same for both the computational and the Hadamard bases (that is, $\beta^{\mathrm{c}}=\beta^{\mathrm{H}}$) and is characterized by
@@ B.3. A Reversed-Space Attack
$\beta^{\mathrm{H}}_{\begin{array}[t]{l}\scriptstyle k=\{t^{\prime}_{-1},t^{\prime}_{0},t^{\prime}_{1},t^{\prime}_{2}\},\\[-3.0pt] \scriptstyle j=\{s_{0},s_{1},s_{2},d_{0},d_{1},d_{2}\}\end{array}}\!\!\!\!\!\!=\frac{1}{2}\left(\begin{array}[]{cccccc}-1&0&0&i&0&0\\ 1&-1&0&i&i&0\\ 0&1&-1&0&i&i\\ 0&0&1&0&0&i\end{array}\right),$
which is immediately obtained by extending Eq. (25) with $\phi=0$ to times $t^{\prime}_{-1}$ and $t^{\prime}_{2}$.
@@
Denote with $B=\{|{V}\rangle$, $|{d_{0}}\rangle$, $|{d_{1}}\rangle$, $|{d_{2}}\rangle$, $|{s_{0}}\rangle$, $|{s_{1}}\rangle$, $|{s_{2}}\rangle\}$ a basis of Bob’s measured space $\mathcal{H}^{\mathrm{B}}$. When Bob measures in the Hadamard basis, he interprets his measurement in the following way, $J_{0}=\{|{d_{1}}\rangle\}$; $J_{1}=\{|{s_{1}}\rangle\}$; $J_{\operatorname{loss}}=B\setminus(J_{0}\cup J_{1})$; and $J_{\operatorname{invalid}}=\emptyset$. [^7] We limit the analysis to pulses that include at most a single photon. Under this assumption, there are no invalid states for this setting. Consider the case where Alice sends $|{+}\rangle$, namely, $\alpha_{t^{\prime}_{0}}=\alpha_{t^{\prime}_{1}}=\frac{1}{\sqrt{2}}$. An error occurs if Bob measures $|{s_{1}}\rangle$, $J_{\operatorname{error}}=\{|{s_{1}}\rangle\}$, and by Observation 1, our attack is required to satisfy
(28) $\frac{1}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{0}}|{E_{t^{\prime}_{0},t^{\prime}_{0}}}\rangle_{\mathrm{E}}+\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle_{\mathrm{E}})+\frac{1}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle_{\mathrm{E}}+\epsilon_{t^{\prime}_{1},t^{\prime}_{1}}|{E_{t^{\prime}_{1},t^{\prime}_{1}}}\rangle_{\mathrm{E}})=0.$
@@
(29) $\frac{i}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{0}}|{E_{t^{\prime}_{0},t^{\prime}_{0}}}\rangle_{\mathrm{E}}-\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle_{\mathrm{E}})+\frac{i}{2\sqrt{2}}(\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle_{\mathrm{E}}-\epsilon_{t^{\prime}_{1},t^{\prime}_{1}}|{E_{t^{\prime}_{1},t^{\prime}_{1}}}\rangle_{\mathrm{E}})=0\text{.}$
@@
(30) $\begin{split}i\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle+i\epsilon_{t^{\prime}_{0},t^{\prime}_{2}}|{E_{t^{\prime}_{0},t^{\prime}_{2}}}\rangle=0,\quad\quad\quad-\epsilon_{t^{\prime}_{0},t^{\prime}_{1}}|{E_{t^{\prime}_{0},t^{\prime}_{1}}}\rangle+\epsilon_{t^{\prime}_{0},t^{\prime}_{2}}|{E_{t^{\prime}_{0},t^{\prime}_{2}}}\rangle=0,\end{split}$
@@
(31) $\begin{split}i\epsilon_{t^{\prime}_{1},t^{\prime}_{-1}}|{E_{t^{\prime}_{1},t^{\prime}_{-1}}}\rangle+i\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle=0,\quad\quad\quad-\epsilon_{t^{\prime}_{1},t^{\prime}_{-1}}|{E_{t^{\prime}_{1},t^{\prime}_{-1}}}\rangle+\epsilon_{t^{\prime}_{1},t^{\prime}_{0}}|{E_{t^{\prime}_{1},t^{\prime}_{0}}}\rangle=0,\end{split}$
@@
(32) $\begin{split}|{0}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{A}}&\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}p|{\phi}\rangle_{\mathrm{E}}|{t^{\prime}_{0}}\rangle_{\mathrm{P}}+p_{1}|{\phi_{1}}\rangle_{\mathrm{E}}|{t^{\prime}_{-1}}\rangle_{\mathrm{P}}+p_{2}|{\psi_{0}}\rangle_{\mathrm{E}}|{V}\rangle_{\mathrm{P}},\\ |{0}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{A}}&\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}p|{\phi}\rangle_{\mathrm{E}}|{t^{\prime}_{1}}\rangle_{\mathrm{P}}+p_{3}|{\phi_{2}}\rangle_{\mathrm{E}}|{t^{\prime}_{2}}\rangle_{\mathrm{P}}+p_{4}|{\psi_{1}}\rangle_{\mathrm{E}}|{V}\rangle_{\mathrm{P}},\end{split}$
@@
with orthogonal $|{E_{1}}\rangle$, $|{E_{2}}\rangle$. As shown in Section 8, this attack is related to the “faked states” attack family (makarov_faked_2005, ; makarov_faked_2008, ) described in Section 3.2.3.